# Discovering Lie Groups with Flow Matching

### v1

@@ Appendix C Results
Figure 11 shows the time progression of $x_{t}$ when generating $5$ samples over $20$ steps for $\mathrm{GL}(2)^{+}$ to $C_{4}$. The gray arrow shows the original $x_{1}$ and the color represents the determinant of the generated transformation matrix. We can see that the transformed point clouds converge to the closest group element in the orbit.
C.2 3D Datasets: with time sampled from uniform distribution
@@
Figure 12: 3D Datasets: A 2D PCA visualization of trajectories from $t=0$ (left) to $t=1$ (right) of the centroids of the transformed objects over $100$ samples.
Figure 12 visualizes the trajectories of the centroids over time for $100$ samples, and PCA is performed to project them onto the 2D plane. We can see that our model seems to learn $\mathrm{Tet}$ and $\mathrm{SO}(2)$ elements, creating some visible clusters, but fails to discover $\mathrm{Oct}$ symmetries.
Figure 13: $\mathrm{SO}(3)$ to $\mathrm{Tet}$: generated transformed point clouds from random $x_{1}$ in the test set.
Figure 14: $\mathrm{SO}(3)$ to $\mathrm{SO}(2)$: generated transformed point clouds from random $x_{1}$ in the test set.
Figure 15: $\mathrm{SO}(3)$ to $\mathrm{Oct}$: generated transformed point clouds from random $x_{1}$ in the test set.
Figure 13, Figure 14 and Figure 15 show generated transformed point clouds from random $x_{1}$ in the test set for $\mathrm{Tet}$, $\mathrm{SO}(2)$ and $\mathrm{Oct}$ respectively. We can see several repeated point clouds, and see that some samples are rotations or reflections of other samples.
@@
We also show the intermediate $x_{t}$ during generation in Figure 16, Figure 17 and Figure 18. For the $\mathrm{Tet}$, we see that the transformed point clouds converge to the closest group element in the orbit. For the $\mathrm{SO}(2)$ group, we can see that the model produces rotations to turn one node of the tetrahedron to be aligned with the $z$-axis and one triangle face to be in the $xy$-plane, which corresponding to $\mathrm{SO}(2)$ symmetries around the $z$-axis. For the $\mathrm{Oct}$ group, we can see that the FM model produces outputs close to the identity at most steps and fails to properly discover $\mathrm{Oct}$.
C.3 3D Datasets: with time sampled from power distribution
In this section, we show additional results for the 3D datasets with time sampled from the power distribution. The power distribution density is shown in Figure 19.
Figure 19: Power distribution density
(a) $\mathrm{SO}(3)$ to $\mathrm{Oct}$
(b) $\mathrm{SO}(3)$ to $\mathrm{Ico}$
@@
Figure 20: 3D datasets with power time schedule: visualization of 5,0005{,}000 generated elements of $\mathrm{SO}(3)$ by converting them to Euler angles. The first two angles are represented spatially on the sphere using Mollweide projection and the color represents the third angle. The elements are canonicalized by the original random transformation and the ground truth elements of the target group are shown in circles with black borders. The points are jittered with uniformly random noise to prevent overlapping.
Figure 20 shows the generated transformations when using the power distribution time schedule. We can see that flow matching now works for octahedral group, but still fails to discover the icosahedral group.
(a) $\mathrm{SO}(3)$ to $\mathrm{Oct}$
(b) $\mathrm{SO}(3)$ to $\mathrm{Ico}$
Figure 21: 3D datasets with power time schedule: A 2D PCA visualization of trajectories from $t=0$ (left) to $t=1$ (right) of the centroids of the transformed objects over $100$ samples.
Figure 21 visualizes the centroid trajectories after performing PCA. The octahedral case shows clear clusters, while the $\mathrm{Ico}$ case seems to output only identity transformations.
C.4 Discovering symmetry in multi-object dataset
In this section, we show additional results for discovering symmetries in a multi-object dataset. We create a dataset consist of four different irregular 3D objects (Figure 22): tetrahedron, triangular prism, cube, octahedron. We use Tet symmetries as the target group and $\mathrm{SO}(3)$ as the prior group. The training details are the same as in the single-object 3D experiments.
Figure 22: Four different irregular 3D objects used in the multi-object dataset
Figure 24: Visualization of 5,0005{,}000 generated elements of $\mathrm{SO}(3)$ by converting them to Euler angles. Same as Figure 20.
Figure 25: Time progression of $x_{t}$ when generating $5$ samples over $100$ steps. The transparent object indicates the original $x_{1}$.
Figure 23 visualizes the centroid trajectories after performing PCA. We can see that the model learns to create clusters corresponding to Tet symmetries. Figure 24 shows the generated transformations, and we can see that the model discovers the Tet group elements. Figure 25 shows the time progression of $x_{t}$ when generating $5$ samples over $100$ steps, and we can see that the transformed point clouds converge to the closest group element in the orbit. Finally, to quantify the symmetry discovery performance, we compute the Wasserstein-1 distance between the discovered group elements and the ground-truth Tet elements, obtaining a distance of $0.10$, while for LieGAN yang2023generative baseline we obtain a distance of 2.202.20(1 Lie generators) 1.951.95(3 Lie generators), and 1.581.58(6 Lie generators), which shows that our method significantly outperforms the baseline in discovering the correct symmetries.
C.5 Discovering structured-preserving transformation
To demonstrate that our method can discover structure-preserving transformations, we create a dataset by applying $z$ axis rotations from $\mathrm{SO}(2)$ with angles sampled from a Gaussian distribution $\mathcal{N}(0,\pi/4)$ to an irregular tetrahedron. Thus, the data distribution is not uniform along the $\mathrm{SO}(2)$ orbits, but the underlying symmetry structure is still $\mathrm{SO}(2)$. We use $\mathrm{SO}(3)$ as the prior group and train our LieFlow model as before.
The results are shown in Figure 26, Figure 27, and Figure 28. Figure 26 visualizes the centroid trajectories after performing PCA. We can see that the model learns to create an incomplete circular pattern, corresponding to leanring a distribution over $\mathrm{SO}(2)$ group elements. Figure 27 shows the generated transformations, and we can see that the elements are concentrated along a one-dimensional manifold, and the samples in the center are more dense than those in the periphery, corresponding to learning a distribution over $\mathrm{SO}(2)$ group elements. Figure 28 shows the histogram of the angles rotated around the $z$-axis extracted from the generated transformations. We can see that the distribution of angles closely matches Gaussian distribution. The Wasserstein-1 distance between the discovered distribution and the ground-truth Gaussian distribution is $0.282$, showing that our method can accurately recover the underlying structure-preserving transformations.
Figure 27: Visualization of 5,0005{,}000 generated elements of $\mathrm{SO}(3)$ by converting them to Euler angles. Same as Figure 20.
Figure 28: Histogram of 5,0005{,}000 angles rotated around the $z$-axis extracted from the generated transformations.
C.6 Flow Matching on Group Elements
This section contains additional visualizations for the flow matching experiment on group elements from $\mathrm{SO}(2)$ to $C_{4}$ as described in Section 5.4.
Figure 29 shows the time progression of $x_{t}$ when generating $100$ samples from $t=0$ to $t=1$. We can see that the samples converge to the closest group element in the orbit.
Figure 30(b) in Figure 30 shows the posterior $p(x_{1}\mid x_{t})$ for all $4$ classes, depending on the original $x_{0}$ and its nearest mode in $\{-\pi,-\pi/2,0,\pi/2\}$. We can see that for all $4$ classes, the posterior $p(x_{1}\mid x_{t})$ remains nearly uniform when $t$ closes to $1$.
(a) $H_{p}(x_{1}\mid x_{t})$
(b) $p(x_{1}\mid x_{t})$
Figure 30 visualizes the posterior $p(x_{1}\mid x_{t})$, divided into 4 buckets depending on the initial position $x_{0}$. We see that the posterior remains nearly uniform until $t=0.95$, and then $x_{t}$ converges to the mode that was the closest when $t=0$.
Figure 31 visualizes how the velocities change over time. Until $t=0.6$, the velocities are close to 0 but slightly push the $x_{t}$’s towards the midpoints between modes. From $t=0.8$ onwards, we see that points $x_{t}$ are pushed to the closest $x_{1}$ mode, with increasing velocities.
Appendix D Flow Matching Directly on Group Elements
Computation of the posterior $p(x_{1}\mid x_{t})$.
To compute the posterior, we use Bayes’ rule as
$p(x_{1}|x_{t})=\frac{p(x_{t}|x_{1})p(x_{1})}{p_{t}(x_{t})},$ (10)
We first sample initial points $x_{0}$ from a uniform prior over $\pi,pi)$ and forward simulate using the learned velocity field $v_{\theta}$ to obtain the trajectories $x_{t}$ and their log probabilities through the continuity equation. To evaluate the likelihood $p(x_{t}\mid x_{1})$ for each candidate $x_{1}$, we invert the linear interpolation $x_{t}=(1-t)x_{0}+tx_{1}$ to find $x_{0}=(x_{t}-tx_{1})/(1-t)$, then integrate the divergence along the trajectory from this inverted $x_{0}$. Special handling is required at the boundaries: at $t=0$, the posterior equals the prior $p(x_{1})$ (which we know in this simple scenario) due to independence, while near $t=1$, we use a Gaussian approximation to avoid numerical instability from division by (1t)(1-t).
Algorithm 4 Compute Posterior $p(x_{1}\mid x_{t})$ for Flow Matching
1:Velocity field $v_{\theta}(x,t)$, prior $p_{0}\sim\mathcal{U}[-\pi,\pi)$, target $p_{1}(x_{1})=\sum_{k=1}^{K}\frac{1}{K}\delta(x_{1}-\mu^{(k)})$, time steps $T$
2:Initialize: Sample $\{x_{0}\}^{N}\sim p_{0}$
3:
4:Forward Simulation:
5:for $t\in\{0,\frac{1}{T},\ldots,1\}$ do
6:  $x_{t}\leftarrow\text{ODESolve}(v_{\theta},x_{0},[0,t])$ $\triangleright$ Integrate velocity field
7:  $\log p_{t}(x_{t})\leftarrow\log p_{0}(x_{0})-\int_{0}^{t}\nabla\cdot v_{\theta}(x_{s},s)ds$ $\triangleright$ Continuity equation
8:end for
9:
10:Posterior Computation:
11:for $t\in\{0,\frac{1}{T},\ldots,1\}$ do
12:  if $t=0$ then
13:   $p(x_{1}\mid x_{0})\leftarrow p_{1}(x_{1})$ $\triangleright$ Independent at $t=0$
14:  else if $t\approx 1$ then
15:   for $k=1$ to $K$ do
16:     $\log p(x_{t}\mid x_{1}^{(k)})\leftarrow-\frac{\|x_{t}-x_{1}^{(k)}\|^{2}}{2\sigma^{2}}$, where $\sigma=0.1$ $\triangleright$ Gaussian Approximation
17:   end for
18:  else
19:   for $k=1$ to $K$ do
20:     $x_{0}^{(k)}\leftarrow\frac{x_{t}-t\cdot x_{1}^{(k)}}{1-t}$ $\triangleright$ Invert linear interpolation
21:     Forward simulate path from $x_{0}^{(k)}$ to time $t$ using $v_{\theta}$ and continuity equation
22:     $\log p(x_{t}\mid x_{1}^{(k)})\leftarrow\log p_{0}(x_{0}^{(k)})-\int_{0}^{t}\nabla\cdot v_{\theta}(x_{s},s)ds$
23:   end for
24:  end if
25:  for $k=1$ to $K$ do
26:   $\log p(x_{1}^{(k)}\mid x_{t})\leftarrow\log p(x_{t}\mid x_{1}^{(k)})+\log p_{1}(x_{1}^{(k)})$ $\triangleright$ Bayes’ Rule
27:  end for
28:
29:  $p(x_{1}^{(k)}\mid x_{t})\leftarrow\frac{\exp(\log p(x_{1}^{(k)}\mid x_{t}))}{\sum_{k}\exp(\log p(x_{1}^{(k)}\mid x_{t}))}$ $\triangleright$ Normalize over $K$ components
30:end for
31:return Posterior $\{p(x_{1}\mid x_{t})\}$ for all time steps

### v2

@@ Appendix F Additional Results
Figures 16 shows the time progression of $x_{t}$ when generating $5$ samples over $20$ steps for $\mathrm{GL}(2)^{+}$ to $C_{4}$. The gray arrow shows the original $x_{1}$ and the color represents the determinant of the generated transformation matrix. We can see that the transformed point clouds converge to the closest group element in the orbit.
F.2 3D irregular tetrahedron
@@
(d) $\mathrm{SO}(3)$ to $\mathrm{Ico}$ (with power time schedule)
Figures 17 visualizes the trajectories of the centroids over time for $100$ samples, and PCA is performed to project them onto the 2D plane. We can see that our model seems to learn $\mathrm{Tet}$, $\mathrm{Oct}$ $\mathrm{Ico}$ and $\mathrm{SO}(2)$ elements, creating some visible clusters.
@@
Figure 20: $\mathrm{SO}(3)$ to $\mathrm{Ico}$: time progression of $x_{t}$ when generating $5$ samples over $100$ steps. The transparent tetrahedron indicates the original $x_{1}$. Training with power time schedule.
@@
We also show the intermediate $x_{t}$ during generation in Figure 18, Figure 19, Figure 20 and Figure 21 for $\mathrm{Tet}$, $\mathrm{Oct}$, $\mathrm{Ico}$ and $\mathrm{SO}(2)$ respectively. For the $\mathrm{Tet}$, $\mathrm{Oct}$, and $\mathrm{Ico}$, we see that the transformed point clouds converge to the closest group element in the orbit instead of the original $x_{1}$, which shows that the model has learned the symmetry structure of the data. For the $\mathrm{SO}(2)$ group, we can see that the model produces rotations to turn one node of the tetrahedron to be aligned with the $z$-axis and one triangle face to be in the $xy$-plane, which corresponds to $\mathrm{SO}(2)$ symmetries around the $z$-axis.
F.3 ModelNet10
Visualization results for rotated ModelNet10 are shown in Figures 22. The generated group elements are clustered around the ground truth group elements for all target groups, which means that our model successfully discovered the symmetries in the rotated ModelNet10 dataset. Note that for the $\mathrm{Ico}$ group, the situation here is even worse than the irregular tetrahedron case without power time schedule, as the model fails to discover any symmetry due to the increased complexity of the dataset. However, with the power time schedule, the model still is able to discover the $\mathrm{Ico}$ symmetries, even with some points masked during training. This shows the effectiveness and robustness of our model combined with the power time schedule in discovering complex symmetries in real-world datasets.
(a) $\mathrm{SO}(3)$ to $\mathrm{Tet}$
(b) $\mathrm{SO}(3)$ to $\mathrm{Oct}$
(c) $\mathrm{SO}(3)$ to $\mathrm{SO}(2)$ over z axis
(d) $\mathrm{SO}(3)$ to $\mathrm{SO}(2)$ over $(0,\frac{1}{2},-\frac{\sqrt{3}}{2})$ axis
(e) $\mathrm{SO}(3)$ to $\mathrm{Ico}$, without power schedule
(f) $\mathrm{SO}(3)$ to $\mathrm{Ico}$, with power schedule
(g) $\mathrm{SO}(3)$ to $\mathrm{Ico}$, with power schedule, 5% masked points
(h) $\mathrm{SO}(3)$ to $\mathrm{Ico}$, with power schedule, 10% masked points
@@ F.3 ModelNet10
Figure 22: Modelnet10 results: Visualization of 5,0005{,}000 generated elements of $\mathrm{SO}(3)$ by converting them to Euler angles. The first two angles are represented spatially on the sphere using Mollweide projection and the color represents the third angle. The elements are canonicalized by the original random transformation and the ground truth elements of the target group are shown in circles with gray borders. The points are jittered with uniformly random noise to prevent overlapping.